Triple Strike, owned and managed by Mr. Brian, showed owner’s equity as of January 1, 2019 as P120,000. It is estimated that at December 31, 2019, total assets will be P400,000 and total liabilities P200,000. How much should be the 2019 projected profit or loss given the following additional information? Each case is to be considered as independent from the others:

     a. Brian did not make any additional investment or withdrawal.

     b. Brian made a cash withdrawal of P50,000 but no additional investment.

     c.  Brian made an additional investment of P125,000 but no cash withdrawal.

     d. Brian made an additional investment of P125,000 and a cash withdrawal of P50,000.



1
Expert's answer
2021-02-22T13:59:48-0500

a) No additional investment and no withdrawal

Total asset + total liabilities = Shareholders' equity + retained profits

400,000+200,000 = 120,000 + X

X=600,000-120,000

=P480,000

b) Cash withdrawal of P50,000 and no additional investment

Withdrawals reduces the asset level

Total asset- cash withdrawal+ total liabilities = Shareholders' equity + retained profits

400,000-50,000+200,000 = 120,000 + X

X=550,000-120,000

=P430,000

c) Additional investment of P125,000 and no withdrawal

Additional investment increase the amount of total assets

Total asset+ additional investment+ total liabilities = Shareholders' equity + retained profits

400,000+125,000+200,000 = 120,000 + X

X=725,000-120,000

=P605,000

d) Cash withdrawal of P50,000 and additional investment of P125,000

Investment increases asset while withdrawal reduce the cash (current assets)

Total asset + additional investment - cash withdrawal+ total liabilities = Shareholders' equity + retained profits

400,000+125,000-50,000+200,000 = 120,000 + X

X=675,000-120,000

=P555,000